pamela broke the fan belt on her car. the belt goes around two pulleys whose centers are 15 inches apart and each pulley is 4 inches in diameter. how long should the belt be?
Answer by mananth(16949)

the perimeter of the wheel = pi*d
3.14*4=12.56
the distance between the centers is 15 inches
2*15 + 12.56=42.56 inches
We take the perimeter of 1 wheel because the belt covers half perimeter on each wheel.
2 times the distance between centers because it loops around